Circular No. 5719

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1985S IN IRAS 0225–103 W. C. Keel, Department of Physics and Astronomy, University of Al-

abama, reports his discovery of an apparent supernova in IRAS 0225–103 (= MCG –2-07-010 = Markarian 1039; α = 2^{h}25^{m}.1, δ = −10^{o}23′, equinox 1950.0). The discovery was made during an analysis of spectroscopic observations (range 330-730 nm) made on 1985 Sept. 19 UT with the 2.5-m Isaac Newton Telescope by M. H. K. de Grijp. Other participants in the project were G. Miley and J. Lub. The rough magnitude was V ∼ 12.8 (V flux = 3.0 × 10^{−14} erg cm^{−2} s^{−1} A^{−1}). SN 1985S is located 15′′ southeast of the galaxy’s nucleus. No bright object appears in this area on the Markarian finding chart (taken from the POSS O print; Markarian et al. 1977, Astrophys. 13, 221). The spectrum, published by de Grijp et al. (1992, A.Ap. Suppl. 96, 389), resembles spectra of certain type-II supernovae, such as that of SN 1988A in NGC 4579 as seen 20 days after maximum light and that of SN 1987A between 60 and 80 days after outburst.

VARIABLE OBJECT NEAR NGC 4192 C. Pollas, Observatoire de la Côte d’Azur, reports that the information

he provided on IAUC 5499 is to be disregarded. He did find an object of V ∼ 19 that is not obviously variable on OCA photographs taken in 1983, 1991, and 1992. The object is also present on the Palomar Sky Survey prints (blue mag ∼ 20, red mag ∼ 18.5). Pollas provides the following accurate position: α = 12^{h}11^{m}16^{s}.39, δ = +15^{o}11′03′′.7 (equinox 1950.0), which corresponds to ∼ 17′′.8 east and 22′′.2 north (± 1′′.5) of the galaxy’s diffuse center.

SUPERNOVA 1993G IN NGC 3690 G. Forti, Osservatorio Astrofisico Arcetri, reports the following po-

sition for SN 1993G as measured by A. Boattini and M. Tombelli: α = 11^{h}25^{m}43^{s}.99, δ = +58^{o}50′02′′.7 (equinox 1950.0); the corresponding offset from the galaxy’s nucleus is 0′′.7 west and 16′′.3 south.

W. Herbst and G. Vinton report a measurement of V = 16.7 ± 0.1 for SN 1993G on Mar. 10.16 UT, using the 0.6-m Perkin reflector (+ CCD) of Van Vleck Observatory, Wesleyan University. A star ∼ 0′.5 east and 2′.7 south of the supernova has V = 11.4 ± 0.1.

1993 March 11 Daniel W. E. Green